# Computing Evolutionarily Stable Strategies in Multiplayer Games

### v1

@@ 1 Introduction
Many evolutionary models of tumor ecology involve frequency-dependent interactions among three or more cancer cell phenotypes (e.g., Basanta et al. (2008, 2011); Archetti (2013); Świerniak and Krześlak (2016)). These systems are naturally formulated as symmetric $n$-player games, and ESS provides a biologically meaningful stability concept. Motivated by this, we develop an algorithm for computing evolutionarily stable strategies in multiplayer symmetric normal-form games. Evolutionarily stable strategies in multiplayer normal-form games have been previously studied and characterized in the biological and mathematical literature (e.g., Peña et al. (2014); Peña and Nöldeke (2015); Gokhale and Traulsen (2010); Wu et al. (2013)).
@@
$u_{i}(\mathbf{x}^{(1)},\ldots,\mathbf{x}^{(n)})=\sum_{s_{1},\ldots,s_{n}\in S}\mathbf{x}^{(1)}_{s_{1}},\ldots,\mathbf{x}^{(n)}_{s_{n}}(u_{i}(s_{1},\ldots,s_{n})).$
@@
$u_{i}(\mathbf{x})=u_{i}(\mathbf{x},\ldots,\mathbf{x})=\sum_{s_{1},\ldots,s_{n}\in S}x_{s_{1}},\ldots,x_{s_{n}}(u_{i}(s_{1},\ldots,s_{n})).$
@@ 2 Algorithm
We first check whether $\mathbf{x}$ is a strict Nash equilibrium. It is well known that in symmetric games any strict SNE is an ESS Maynard Smith and Price (1973); Maynard Smith (1982); Hofbauer and Sigmund (1998). Moreover, among symmetric Nash equilibria, strictness is equivalent to being a pure strategy whose chosen action is the unique best response to itself Maynard Smith (1982); Hofbauer and Sigmund (1998). Consequently, whenever $\mathbf{x}$ is pure and satisfies $|BR(\mathbf{x})|=1$, we may immediately conclude that $\mathbf{x}$ is an ESS. We next test whether a pure mutant can invade $\mathbf{x}.$ After applying these efficient preprocessing tests, we apply Algorithm 4 to determine whether $\mathbf{x}$ can be invaded by any mixed mutant.
@@
6:if $|\mathrm{BR}(\mathbf{x})|=1$ and $\mathbf{x}$ is pure then
@@
18:if $\text{status}=\text{INFEASIBLE}$ then
@@
20:end if
21:if $F^{\star}<-\epsilon_{p}$ then
@@
25:else
[…]
27:end if
Algorithm 4 tests whether an SNE $\mathbf{x}$ can be invaded by any mixed mutant by creating and solving a quadratically-constrained quadratic program (QCQP). The variables $y_{i}$ represent the non-negative strategy probabilities over the elements of $\mathrm{BR}(\mathbf{x}).$ Let $\overline{\mathbf{y}}$ denote the extension of $\mathbf{y}$ to a $K$-dimensional strategy (by setting entries not in $\mathrm{BR}(\mathbf{x})$ to 0). In order for $\mathbf{x}$ to be an ESS, we must ensure that $U(\mathbf{x},\overline{\mathbf{y}},\mathbf{x})>U(\overline{\mathbf{y}},\overline{\mathbf{y}},\mathbf{x})$ for all mixed strategies $\overline{\mathbf{y}}\neq\mathbf{x}.$ For simplicity of notation we use $\mathbf{y}$ instead of defining $\overline{\mathbf{y}}$ in the pseudocode. Numerically, we say that $\overline{\mathbf{y}}\neq\mathbf{x}$ when $\overline{\mathbf{y}}-\mathbf{x}||_{2}\geq\delta$.
@@
11:if QCP is infeasible then
12:  return false $\triangleright$ should not occur if $\mathbf{x}$ was feasible
13:end if
14: Let $D^{\star}$ denote the optimal objective value.
15:if $D^{\star}>\epsilon_{\text{dist}}$ then
16:  return true $\triangleright$ another SNE exists on $\mathbf{T}$: degenerate
17:else
18:  return false $\triangleright$ unique SNE on $\mathbf{T}$ (nondegenerate)
19:end if
@@ 3 Experiments
2 2 All mixed strategies supported on 0,1 (continuum) $\varnothing$
@@ 4 Conclusion
While our algorithm is only guaranteed to find all ESSs in nondegenerate games, it is still guaranteed to find a subset of all ESSs if run on degenerate games. Thus, the algorithm may still be useful in practice even on degenerate games. Three of the eight example games considered are degenerate; on the first the algorithm correctly classified that no ESSs exist, on the second the algorithm correctly found the unique ESS, and on the third the algorithm found three of the infinitely many ESSs. Future work can explore improvements for solving degenerate games. For example, once an SNE $\mathbf{x}$ is found for given support $\mathbf{T}$ (by solving the QCP in Algorithm 2), we can then solve a second QCP that searches for an additional SNE $\mathbf{x}^{\prime}$ that satisfies the same constraints with objective of maximizing the Euclidean distance between $\mathbf{x}^{\prime}$ and $\mathbf{x}.$ We can continue this procedure indefinitely to keep finding additional SNE which are ESS candidates until the objective value falls below a numerical tolerance. This may be helpful for specific games of interest which are known to be degenerate (e.g., by applying the degeneracy-testing procedure in Algorithm 5), though it will not be able to solve the general problem of finding all SNE, since the set of SNE for a given support may be infinite (as several of the example games indicate).
@@ Appendix A Payoff Tensors for the Example Games
$A[0]=\begin{pmatrix}0&-1\\ 2&0\end{pmatrix},\qquad A[1]=\begin{pmatrix}2&0\\ -1&0\end{pmatrix}$.
@@
$A[0]=\begin{pmatrix}2&0&0\\ 0&0&0\\ 0&0&0\end{pmatrix},\qquad A[1]=\mathbf{0},\qquad A[2]=\mathbf{0}.$
@@
$A[0]=\begin{pmatrix}2.0&1.0&0.5\\ 3.0&2.5&1.0\\ 1.5&1.0&0.5\end{pmatrix},\quad A[1]=\begin{pmatrix}1.5&1.0&0.5\\ 1.0&0.8&0.3\\ 0.8&0.5&0.2\end{pmatrix},\quad A[2]=\begin{pmatrix}2.0&1.5&1.0\\ 1.5&1.0&0.8\\ 1.0&0.8&0.5\end{pmatrix}$.

### v4

# Computing ESS in Multiplayer Games

@@ 1 Introduction
Many evolutionary models of tumor ecology involve frequency-dependent interactions among three or more cancer cell phenotypes (e.g., Basanta et al. (2008, 2011); Archetti (2013); Swierniak and Krzesłak (2016)). These systems are naturally formulated as symmetric $n$-player games, and ESS provides a biologically meaningful stability concept. Motivated by this, we develop an algorithm for computing evolutionarily stable strategies in multiplayer symmetric normal-form games. Evolutionarily stable strategies in multiplayer normal-form games have been previously studied and characterized in the biological and mathematical literature (e.g., Peña et al. (2014); Peña and Nöldeke (2015); Gokhale and Traulsen (2010); Wu et al. (2013)).
@@
$u_{i}(\mathbf{x}^{(1)},\ldots,\mathbf{x}^{(n)})=\sum_{s_{1},\ldots,s_{n}\in S}x^{(1)}_{s_{1}}\cdots x^{(n)}_{s_{n}}u_{i}(s_{1},\ldots,s_{n}).$
@@
$u_{i}(\mathbf{x})=u_{i}(\mathbf{x},\ldots,\mathbf{x})=\sum_{s_{1},\ldots,s_{n}\in S}x_{s_{1}}\cdots x_{s_{n}}u_{i}(s_{1},\ldots,s_{n}).$
@@ 2 Algorithm
We first check whether $\mathbf{x}$ is a strict Nash equilibrium. It is well known that in symmetric games any strict SNE is an ESS Maynard Smith and Price (1973); Maynard Smith (1982); Hofbauer and Sigmund (1998). Moreover, among symmetric Nash equilibria, strictness is equivalent to being a pure strategy whose chosen action is the unique best response to itself Maynard Smith (1982); Hofbauer and Sigmund (1998). Consequently, whenever $\mathbf{x}$ is pure and satisfies $|BR(\mathbf{x})|=1$, we may immediately conclude that $\mathbf{x}$ is an ESS. Note that $|BR(\mathbf{x})|=1$ implies that $\mathbf{x}$ is pure, so we can simplify this procedure to just check if $|BR(\mathbf{x})|=1.$ We next test whether a pure mutant can invade $\mathbf{x}$ by iterating over all pure strategies $i\in\mathrm{BR}(\mathbf{x})$ and testing whether
$U(\mathbf{e}_{i},\mathbf{e}_{i},\mathbf{x})\geq U(\mathbf{x},\mathbf{e}_{i},\mathbf{x})+\epsilon_{p}.$
After applying these efficient tests, we apply Algorithm 4 to determine whether $\mathbf{x}$ can be invaded by any mixed mutant.
@@
6:if $|\mathrm{BR}(\mathbf{x})|=1$ then
@@
18:if $F^{\star}<-\epsilon_{p}$ then
@@
20:else if $F^{\star}>\epsilon_{p}$ then
@@
24:end if
Algorithm 4 tests whether an SNE $\mathbf{x}$ can be invaded by any mixed mutant by creating and solving a quadratically-constrained quadratic program (QCQP). The variables $y_{i}$ represent the non-negative strategy probabilities over the elements of $\mathrm{BR}(\mathbf{x}).$ Let $\overline{\mathbf{y}}$ denote the extension of $\mathbf{y}$ to a $K$-dimensional strategy (by setting entries not in $\mathrm{BR}(\mathbf{x})$ to 0). In order for $\mathbf{x}$ to be an ESS, we must ensure that $U(\mathbf{x},\overline{\mathbf{y}},\mathbf{x})>U(\overline{\mathbf{y}},\overline{\mathbf{y}},\mathbf{x})$ for all mixed strategies $\overline{\mathbf{y}}\neq\mathbf{x}.$ For simplicity of notation we use $\mathbf{y}$ instead of defining $\overline{\mathbf{y}}$ in the pseudocode. Numerically, we say that $\overline{\mathbf{y}}\neq\mathbf{x}$ when $\overline{\mathbf{y}}-\mathbf{x}||_{2}\geq\delta$. The QCQP is always feasible since we must have $|BR(\mathbf{x})|\geq 2$.
@@
11: Let $D^{\star}$ denote the optimal objective value.
12:if $D^{\star}>\epsilon_{\text{dist}}$ then
13:  return true $\triangleright$ another SNE exists on $\mathbf{T}$: degenerate
14:else
15:  return false $\triangleright$ unique SNE on $\mathbf{T}$: nondegenerate
16:end if
@@ 3 Experiments
2 2 All mixed strategies $\varnothing$
@@ 4 Conclusion
While our algorithm is only guaranteed to find all ESSs in nondegenerate games, it is still guaranteed to find a subset of all ESSs if run on degenerate games. Thus, the algorithm may still be useful in practice even on degenerate games. Two of the eight example games considered are degenerate; on the first the algorithm correctly classified that no ESSs exist, and on the second the algorithm correctly found the unique ESS. Future work can explore improvements for solving degenerate games. For example, once an SNE $\mathbf{x}$ is found for given support $\mathbf{T}$ (by solving the QCP in Algorithm 2), we can then solve a second QCP that searches for an additional SNE $\mathbf{x}^{\prime}$ that satisfies the same constraints with objective of maximizing the Euclidean distance between $\mathbf{x}^{\prime}$ and $\mathbf{x}.$ We can continue this procedure indefinitely to keep finding additional SNE which are ESS candidates until the objective value falls below a numerical tolerance. This may be helpful for specific games of interest which are known to be degenerate (e.g., by applying the degeneracy-testing procedure in Algorithm 5), though it will not be able to solve the general problem of finding all SNE, since the set of SNE for a given support may be infinite (as several of the example games indicate).
@@ Appendix A Payoff Tensors for the Example Games
$A[0]=\begin{pmatrix}-3&0\\ 0&-3\end{pmatrix},\qquad A[1]=\begin{pmatrix}1&-3\\ -3&-1\end{pmatrix}$.
@@
$A[0]=\begin{pmatrix}2&0&0\\ 0&0&0\\ 0&0&0\end{pmatrix},\qquad A[1]=\begin{pmatrix}0&0&0\\ 0&0&0\\ 0&0&0\end{pmatrix},\quad A[2]=\begin{pmatrix}0&0&0\\ 0&0&0\\ 0&0&0\end{pmatrix}$.
@@
$A[0]=\begin{pmatrix}4&2&3\\ 2&0&1\\ 3&1&2\end{pmatrix},\quad A[1]=\begin{pmatrix}0&0&0\\ 0&0&0\\ 0&0&0\end{pmatrix},\quad A[2]=\begin{pmatrix}2&1&3\\ 1&0&2\\ 3&2&4\end{pmatrix}$.